a student takes a 8- question, multiple-choice exam with 2 choices for each question and guesses on each question. find the probability of guessing at least 3 out of 8 correctly
Answer by stanbon(75887)

a student takes a 8- question, multiple-choice exam with 2 choices for each question and guesses on each question. find the probability of guessing at least 3 out of 8 correctly
----
Binomial Problem with n = 8 and p(correct) = 1/2
-----
P(x=3) = 8C3*(1/20^3*(1/2)^5 = 56/2^8 = 0.2188
==============================
Cheers,
Stan H.